# The origin of anomalous non-linear microwave absorption in Josephson junction qubits: mysterious nature of two level systems or their dynamic interaction?

### v1

@@ III.1 TLSs with distributed dipole moments: Intensity dependence of absorption
Here we calculate the loss-tangent for TLSs with distributed dipole moments. We assume that TLS dipole moments $p$ are distributed between their minimum and maximum values $p_{\rm min}$ and $p_{\rm max}$, respectively, such that $p_{\rm min}\ll p_{\rm max}$ and introduce their hypothetical distribution in the form
$P(\Delta_{0},\Delta,p)=\frac{P_{0}}{\Delta_{0}}F(p),~{}F(p)\approx\frac{p_{min}^{2-\eta}}{(2-\eta)p^{3-\eta}},$
$p_{\rm min}<p<p_{\rm max},~{}0\leq\eta<1.$ (20)
@@
Assume that there is no correlations between TLS elastic tensors and dipole moments and they are coupled approximately identically to the elastic field, which is a standard assumption for the TLS model. Then resonant TLSs relaxation times are connected to their minimum relaxation time $T_{\rm 1}$ as $T_{1}(\Delta_{0})=T_{\rm 1}(E/\Delta_{0})^{2}$ [33; 41]. TLS coupling to the external field is determined by its Rabi frequency that can be expressed in terms of the maximum Rabi frequency $\Omega_{max}=Fp_{max}$ (remember that $F$ stands for the external microwave field) as $\Omega=(\Delta_{0}/E_{0})\Omega_{max}\cos(\theta)p/p_{max}$, where $\theta$ is the angle between the field and the TLS dipole moment (remember that for resonant TLSs $E_{0}\approx\hbar\omega$).
@@
The loss tangent is defined by the average sum of all TLS contributions within the unit volume, which is equivalent to its averaging over TLS tunneling amplitudes $\Delta_{0}$, angles $\theta$ and TLS dipole moments absolute values $p$ with the probability density function Eq. (20). The integral can be conveniently expressed as
@@
There are three distinguishable behaviors of the loss tangent depending on the product of the maximum Rabi frequency and the minimum relaxation time. For very small power, $\Omega_{\rm max}T_{\rm min}\ll 1$ the denominator in the integrand in Eq. (21) can be replaced with $1$ and the loss tangent is intensity independent approaching its maximum ($\tan(\delta_{\rm lin})$, cf. Eq. (4)). At intermediate powers $\Omega_{\rm min}<1/T_{\rm min}<\Omega_{\rm max}$ ($\Omega_{\rm min}=\Omega_{\rm max}R$) the main contribution to the absorption is originated from dipole moments $p\sim p_{\rm nl}$ where the saturation of absorption begins, i. e. $p_{\rm nl}F\approx\hbar/T_{\rm min}$. Here the loss tangent is reduced by the volume factor $(p_{\rm nl}/p_{\rm max})^{3}$ and enhanced by the probability density contribution $(p_{\rm nl}/p_{\rm max})^{-3+\eta}$. Then the loss tangent depends on the external field as $\tan(\delta)\propto p_{\rm nl}^{\eta}\propto F^{-\eta}$. For $\eta=0$ this dependence becomes logarithmic in the form $\tan(\delta)\propto\ln(\hbar/(T_{1,min}Fp_{\rm min})$, that is exactly the dependence that was targeted in Ref. [17]. Finally for $1/T_{\rm min}<\Omega_{\rm max}R$ the inverse field dependence of loss tangent takes place as in the standard TLS model with all identical dipole moments [21].
@@
We evaluated integrals in Eq. (21) in all these three limits setting the lower and upper integration limits over dipole moments ($z$) to zero and infinity respectively, for intermediate intensities that is justified in case of $10<\eta<$ 1. Then we get
@@
$\tan(\delta)\approx\frac{4\pi^{2}}{3(2-\eta)R^{\eta}}\frac{P_{0}p_{min}^{2}}{\epsilon}\begin{cases}\frac{1-R^{\eta}}{\eta},~{}\Omega_{\rm max}<\frac{1}{T_{\rm min}},\\ \frac{3}{4(3-\eta)}\frac{\Gamma(\eta/2)\Gamma((1-\eta)/2)\Gamma(1+\eta/2)}{\Gamma((3+\eta)/2)}\frac{1}{(\Omega_{\rm max}T_{\rm 1min})^{\eta}},~{}\Omega_{\rm min}<\frac{1}{T_{\rm min}}<\Omega_{\rm max},\\ \frac{3\pi}{8\sqrt{2}R^{1-\eta}(1-\eta)}\frac{1}{\Omega_{\rm max}T_{\rm 1min}},~{}\frac{1}{T_{\rm min}}<\Omega_{\rm min},\\ \end{cases}.$ (22)
@@
where $\Omega_{\rm min}=R\Omega_{\rm max}$ is the typical Rabi frequency for the minimum dipole moment. In the special case of interest of $\eta=0$ the generalization for small and large intensities is quite straightforward. In the intermediate regime of main interest we get (within the logarithmic accuracy)
@@
$\tan(\delta)\approx\frac{\pi^{2}}{3}\frac{P_{0}p_{min}^{2}}{\epsilon}\ln\left(\frac{1}{\Omega_{\rm min}T_{\rm 1min}}\right),~{}\Omega_{\rm min}=\frac{p_{min}F}{\hbar}.$ (23)
@@
Figure 3: Loss tangent dependence on the intensity for the TLS dipole moments distributed as $F(p)\propto p^{-3+\eta}$ for (a) $\eta=0.25$ and (b) $\eta=0.5$. Ratio $R$ of minimum to maximum dipole moments is shown near each graph. Dotted lines show the fit by $1/\Omega_{\rm max}^{\eta}$ and dashed lines show the fit by $1/\Omega_{\rm max}$.
The results of the numerical evaluation of the loss tangent using Eq. (21) are shown in Fig. 3. The obtained dependencies of the loss tangent on the intensity are getting consistent with the analytical asymptotics in Eq. (22) at sufficiently small ratio of minimum to maximum dipole moments. The field dependence gets stronger with increasing the field. A similar trend is seen experimentally (see for example Fig. 5 in Ref. [24] fitted by the logarithmic dependence transferring to the inverse field dependence with increasing field in agreement with Eq. 5 in the same paper). Below we discuss whether the dipole moment power law distribution Eq. (20) is consistent with experimental data where dipole moments of many individual TLSs were measured directly.
III.2 Dipole moment distribution of experimentally observed TLSs
@@
A typical TLS dipole moment can be determined experimentally using a variety of methods. For example two and three pulse dielectric echo measurements permit us to extract a typical value of dipole moment and give some knowledge about its distribution in the integrated form [42; 43; 44]. TLS dipole moment can be estimated measuring absorption in the presence of time-varying bias. However, these methods do not probe the distribution of dipole moments including its tails as expressed by Eq. (20), which are potentially responsible for the anomalous field dependence of the loss tangent.
@@
In all these experiments TLS dipole moments vary by at least a factor of three that is not typical for the narrow Gaussian-like distributions. The histograms expressing dipole moment probability density function were reported in Ref. [39] for poly-crystalline and amorphous aluminum oxide films. In the amorphous sample where the maximum TLS dipole moment is as large as $16$ D the Gaussian distribution does not fit the dipole moment probability density function, while for the poly crystalline sample the maximum dipole moment is around $7$ D and the Gaussian distribution matches data better, yet, not perfect.
Figure 4: The probability density function vs. dipole moment for amorphous Al oxide [39] compared with the theoretical expectations.
To check whether the hypothesis of power law dipole moment distribution Eq. (20) is consistent with the experimental data reported in Ref. [39], we attempted to fit them using the power law probability density function. For the poly-crystalline sample the best power law fit was obtained using $p^{-3+\eta}$ with $\eta\approx-2$ (not shown) and it doesn’t fit the data better than the fit by the Gaussian distribution [39]. However, for the amorphous sample the probability density function follows the $p^{-3}$ law reasonably well as shown in Fig. 4 so it is consistent with the proposed hypothetical distribution.
This analysis is very preliminary and it should hopefully encourage experimentalists to collect more data and verify the relevance of the present model. Below we discuss some theoretical grounds behind the power law dipole moment distribution.
@@
The power law dependence of the probability density function on the dipole moment, $F(p)\propto p^{-3+\eta}$, expressed by Eq. (20), emerges naturally for coupling constants of low energy collective excitations of interacting two-level defects (Ising pseudospins $1/2$) coupled by the long-range interaction $u/r^{d}$ in $d$-dimensions [15; 40]. The quoted theory was developed to interpret the quantitative universality of TLSs in amorphous solids as those collective excitations, following the earlier suggestion of Refs. [13; 47]. The renormalization group analysis of this model with random uncorrelated interaction constants $u$ between each pair of defects [15; 40] predicted that the low energy excitations are distinguished by the number of participating defects $n$ and possess the universal densities of states $P_{n}\propto\ln(n)/n^{2}$. In that specific model the coupling constant for $n$-defect excitations scales with $n$ as $\gamma\propto\sqrt{n}$. The distribution of coupling constants $\gamma$ behaves consequently as
@@
The predicted distribution of dipole moments is relevant if the dipole-dipole interaction is the strongest long-range interaction in the system. To our knowledge in the vast majority of amorphous dielectrics elastic interaction is stronger than the dipole-dipole interaction [49; 16]. Indeed, TLSs of atomic nature possess very small dipole moments of order or less than few D because heavy atoms can tunnel only to very short distances. However, if TLSs are of electronic nature, then the tunneling distances can be much longer up to nanometers and the associated dipole-dipole interaction can exceed the elastic one. Consequently, we expect that the materials showing anomalous non-linear absorption possess TLSs of electronic nature formed by e. g. localized quasi-particles [50; 51].

### v2

@@ III.1 TLSs with distributed dipole moments: Intensity dependence of absorption
Here we calculate the loss-tangent for TLSs with distributed dipole moments. We assume that TLS dipole moments $p$ are distributed between their minimum and maximum values $p_{\rm min}$ and $p_{\rm max}$, respectively, such that $p_{\rm min}\ll p_{\rm max}$ and introduce their hypothetical probability density function in the simple form
$P(\Delta_{0},\Delta,p)=\frac{P_{0}}{\Delta_{0}}F(p),$
$F(p)\approx\begin{cases}0,~{}p<p_{\rm min},\\ \frac{p_{\rm min}^{2-\eta}}{(2-\eta)p^{3-\eta}},~{}p_{\rm min}<p<p_{\rm max},\\ 0,p>p_{\rm max},\end{cases}~{}0\leq\eta<1.$ (20)
This hypothetical function discards the domains $p<p_{\rm min}$ and $p>p_{\rm max}$. It is natural to expect that the function $F(p)$ scales as $F(p)\propto p^{2}$ (cf. Ref. [41]) for $p<p_{\rm min}$ and it rapidly (exponentially or so) decreases for $p>p_{\rm max}$. Under these assumptions the loss tangent intensity dependence is qualitatively correct in various asymptotic regimes considered below with the accuracy to the numerical factor of order of unity, since at high fields the losses will be due to TLSs with $p\sim p_{\rm min}$, while the field independent loss-tangent at small fields is determined by TLSs with dipole moments of comparable to $p_{\rm max}$.
@@
Assume that there is no correlations between TLS elastic tensors and dipole moments and they are coupled approximately identically to the elastic field, which is a standard assumption for the TLS model. Then resonant TLSs relaxation times are connected to their minimum relaxation time $T_{\rm 1,min}$ as $T_{1}(\Delta_{0})=T_{\rm 1,min}(E/\Delta_{0})^{2}$ [33; 42]. Here we use notations different from those in the Introduction and Sec. II, where the same TLS relaxation time $T_{1}$ was used to estimate the fluctuator contribution to the loss tangent, because the absorption is determined by resonant TLSs having $\Delta_{0}\sim E=\hbar\omega$. Consequently, absorbing TLSs possess relaxation times of order of minimum TLS relaxation time denoted there as $T_{1}$. Since here we perform accurate calculations, we assign to each TLS its own relaxation time $T_{1}$ expressed in terms of the minimum TLS relaxation time $T_{\rm 1,min}$ as $T_{\rm 1,min}(E/\Delta_{0})^{2}$.
TLS coupling to the external field is determined by its Rabi frequency that can be expressed in terms of the maximum Rabi frequency $\Omega_{max}=Fp_{max}$ (remember that $F$ stands for the external microwave field) as $\Omega=(\Delta_{0}/E_{0})\Omega_{max}\cos(\theta)p/p_{max}$, where $\theta$ is the angle between the field and the TLS dipole moment (remember that for resonant TLSs $E_{0}\approx\hbar\omega$).
@@
The loss tangent is defined by the average sum of all TLS contributions within the unit volume, which is equivalent to its averaging over TLS tunneling amplitudes $\Delta_{0}$, angles $\theta$ and TLS dipole moments absolute values $p$ with the probability density function, Eq. (20). The integral can be conveniently expressed as
@@
There are three distinguishable behaviors of the loss tangent depending on the product of the maximum Rabi frequency and the minimum relaxation time. For very small power, $\Omega_{\rm max}T_{\rm 1,min}\ll 1$ the denominator in the integrand in Eq. (21) can be replaced with $1$ and the loss tangent is intensity independent approaching its maximum ($\tan(\delta_{\rm lin})$, cf. Eq. (4)). At intermediate powers $\Omega_{\rm min}<1/T_{\rm 1,min}<\Omega_{\rm max}$ ($\Omega_{\rm min}=\Omega_{\rm max}R$) the main contribution to the absorption is originated from dipole moments $p\sim p_{\rm nl}$ where the saturation of absorption begins, i. e. $p_{\rm nl}F\approx\hbar/T_{\rm 1,min}$. Here the loss tangent is reduced by the volume factor $(p_{\rm nl}/p_{\rm max})^{3}$ and enhanced by the probability density contribution $(p_{\rm nl}/p_{\rm max})^{-3+\eta}$. Then the loss tangent depends on the external field as $\tan(\delta)\propto p_{\rm nl}^{\eta}\propto F^{-\eta}$. For $\eta=0$ this dependence becomes logarithmic in the form $\tan(\delta)\propto\ln(\hbar/(T_{1,min}Fp_{\rm min})$, that is exactly the dependence that was targeted in Ref. [17]. Finally for $1/T_{\rm 1,min}<\Omega_{\rm max}R$ the inverse field dependence of loss tangent takes place as in the standard TLS model with all identical dipole moments [21].
@@
We evaluated integrals in Eq. (21) in all these three limits using the probability density function, Eq. (20), and setting the lower and upper integration limits over dipole moments ($z$) to zero and infinity respectively, for intermediate intensities that is justified in case of $10<\eta<$ 1. Then we get
@@
$\tan(\delta)\approx\frac{4\pi^{2}}{3(2-\eta)}\frac{P_{0}p_{min}^{2}}{\epsilon}\begin{cases}\frac{1-R^{\eta}}{R^{\eta}\eta},~{}\Omega_{\rm max}<\frac{1}{T_{\rm 1,min}},\\ \frac{3}{4(3-\eta)}\frac{\Gamma(\eta/2)\Gamma((1-\eta)/2)\Gamma(1+\eta/2)}{\Gamma((3+\eta)/2)}\frac{1}{(\Omega_{\rm min}T_{\rm 1,min})^{\eta}},~{}\Omega_{\rm min}<\frac{1}{T_{\rm 1,min}}<\Omega_{\rm max},\\ \frac{3\pi}{8\sqrt{2}(1-\eta)}\frac{1}{\Omega_{\rm min}T_{\rm 1,min}},~{}\frac{1}{T_{\rm 1,min}}<\Omega_{\rm min},\\ \end{cases}.$ (22)
@@
where $\Omega_{\rm min}=R\Omega_{\rm max}$ is the typical Rabi frequency for the minimum dipole moment. In the special case of interest of $\eta=0$ we get
@@
$\tan(\delta)\approx\frac{2\pi^{2}}{3}\frac{P_{0}p_{min}^{2}}{\epsilon}\begin{cases}\ln(R),~{}\Omega_{\rm max}<\frac{1}{T_{\rm 1,min}},\\ \ln\left(\frac{1}{\Omega_{\rm min}T_{\rm 1,min}}\right),~{}\Omega_{\rm min}<\frac{1}{T_{\rm 1,min}}<\Omega_{\rm max},\\ \frac{3\pi}{8\sqrt{2}}\frac{1}{\Omega_{\rm min}T_{\rm 1,min}},~{}\frac{1}{T_{\rm 1,min}}<\Omega_{\rm min},\\ \end{cases}.$ (23)
The results for small exponents $\eta$, i. e. $\eta<1/|\ln(\Omega_{\rm min}T_{\rm 1,min})|$, given by Eq. (22), are formally identical to the results in Eq. (23) for $\eta=0$. Since both the analysis of experimental data reported in Sec. III.2 and the expectations based on the model of TLS formation due to the long-range interactions reported in Sec. III.3 suggest a nearly vanishing exponent $\eta$ in the power law probability density function of TLS dipole moments, Eq. (20), here we focus on the regime $\eta=0$. We show that the loss tangent field dependence in this regime is consistent with the experimental results of Ref. [24] and, hopefully, with other observations of similar loss tangent behaviors.
@@
Figure 3: Loss tangent dependence on the intensity for the TLS dipole moments distributed as $F(p)\propto p^{-3}$ for $\eta=0$. Ratio $R$ of minimum to maximum dipole moments is shown on the left side of each graph. Dashed lines show the fits by the Rabi frequency dependence in the form $\ln(1+C/(\Omega_{\rm min}\sqrt{T_{\rm 1,min}T_{\rm 2,min}}))$ ($T_{\rm 2,min}=2T_{\rm 1,min}$) used in Ref. [24]. The fitting parameters $C$ are all equal $1.05$ independent of the ratio $R$.
Indeed, experimental data were fitted in Ref. [24] using the loss tangent field dependence in the form $\tan(\delta)\propto\ln(\gamma/\Omega_{R}+C_{1})$ where $\gamma$ is some constant parameter having a frequency dimension, $\Omega_{R}$ is the typical TLS Rabi frequency proportional to the AC field amplitude and $C_{3}\geq 1$ is a constant including the field independent loss of an unknown nature. This loss can be separated from the field dependent contribution to the loss tangent as $\ln(\gamma/\Omega_{R}+C_{1})=\ln(C_{1})+\ln(\gamma_{1}/\Omega_{R}+1)$, where $\gamma_{1}=\gamma/C_{1}$. Our consideration does not include field independent loss, while the field dependence expressed by the logarithmic factor $\ln(\gamma_{1}/\Omega_{R}+1)$ is asymptotically consistent with the predictions of the theory, Eq. (23) obtained using $\gamma_{1}\approx 1/T_{1}$, and $\Omega_{\rm min}$ instead of $\Omega_{R}$. Indeed, we get a logarithmic field dependence $-\ln(\Omega_{\rm min}T_{1})$ at intermediate fields $\Omega_{\rm min}<1/T_{1}<\Omega_{\rm max}$ and inverse field dependence $1/(\Omega_{\rm min}T_{1})$ at large fields $\Omega_{\rm min}>1/T_{1}$, Eq. (23). Moreover, we can almost perfectly fit the theoretical predictions for the loss tangent, Eq. (21), evaluated for $\eta=0$ using the interpolation formula proposed in Ref. [24] as illustrated in Fig. 3. The discrepancy at very small fields corresponds to the linear regime $\Omega_{\rm max}<1/T_{1}$ (first line in Eq. (23)), where the logarithmic interpolation does not work. This regime is not attained in Ref. [24] at classical fields, because the non-linear regime emerges there even at AC fields corresponding to one photon per cavity. Smaller fields cannot be treated classically. The investigation of a quantum regime is beyond the scope of the present work.
The logarithmic interpolation formula was suggested in Ref. [24] based on the prior dynamic interaction based theory [17]. It makes the predictions of the loss tangent field dependence quite similar to those in the present work. As we argued in the Introduction the dynamic interaction based theory underestimated the size of the resonant domain within the non-linear regime, where it is determined by a Rabi frequency proportional to the AC field rather than the TLS decoherence rate. As it is shown in Sec. II, if the definition of a resonant domain would be corrected, then the field dependence is not logarithmic, but it is even stronger than the inverse field dependence realized in the absence of fluctuators. Consequently, we do not believe that the explanation of Ref. [17] is applicable to the experimental data under consideration.
III.2 Comparison of theory with dipole moment probability density function determined by direct measurements [39].
@@
A typical TLS dipole moment can be determined experimentally using a variety of methods. For example two and three pulse dielectric echo measurements permit us to extract a typical value of dipole moment and to obtain some knowledge about its distribution in the integrated form [43; 44; 41]. TLS dipole moment can also be estimated measuring absorption in the presence of time-varying bias. However, these methods do not probe the distribution of dipole moments particularly including its tails as expressed by Eq. (20), which are potentially responsible for the anomalous field dependence of the loss tangent, considered in the present work.
Figure 4: Dipole probability density function vs. dipole moment rescaled by the factor $p^{3}$ for polycrystalline and amorphous Al oxides [39].
@@
[…]
The probability density function was analyzed in Ref. [39]. It was fitted by Gaussian and gamma-distributions. However, this analysis targeted typical TLS dipole moments and it is not relevant for distribution tails, which are the targets of our consideration.
Figure 5: The probability density function vs. dipole moment for amorphous Al oxide [39] compared with the theoretical expectations. The point with zero value of the function is omitted, because it cannot be shown in the logarithmic scale.
The loss tangent is determined by the tail of the probability density function $F(p)$ of dipole moments if it decreases with increasing the dipole moment as $p^{-3}$ or slower. To verify whether this is true or not for the experimental data of Ref. [39] we evaluated the product of that function obtained experimentally and the cube of a dipole moment and report the results in Fig. 4 both for polycrystalline and amorphous aluminum oxide films. According to the graph the distribution function tail is not significant for polycrystalline films, where the probability density function decreases with increasing the dipole moment much faster than $p^{-3}$. However, in the amorphous films the tails are significant and should contribute to the loss tangent.
We also attempted to fit the tails of the probability density functions for dipole moment projections using the power law fit. For the poly-crystalline sample the best power law fit was obtained using a dependence $F(p)\propto p^{-5}$ (not shown). However, for the amorphous sample the probability density function follows the $p^{-3}$ law reasonably well for dipole moments exceeding $4$ D as shown in Fig. 5, so it is consistent with the proposed hypothetical distribution. We fit only the tail of the distribution $p>4D$ assuming that the minimum dipole moment $p_{\rm min}$ is of order of few D. At smaller dipole moments the power law distribution is obviously irrelevant.
Based on the observed behavior of the probability density functions in polycrystalline and amorphous samples we predict that the anomalously weak dependence of loss tangent on the AC field intensity should be seen in amorphous samples, while the non-linear loss tangent in polycrystalline samples should behave according to the standard TLS model. The present work should hopefully encourage experimentalists to collect more data and verify the relevance of the present model. Below we discuss some theoretical grounds behind the power law dipole moment distribution.
@@
The power law dependence of the probability density function on the dipole moment, $F(p)\propto p^{-3+\eta}$, expressed by Eq. (20), emerges naturally for coupling constants of low energy collective excitations of interacting two-level defects (Ising pseudospins $1/2$) coupled by the long-range interaction $u/r^{d}$ in $d$-dimensions [15; 40]. The quoted theory was developed to interpret the quantitative universality of TLSs in amorphous solids as those collective excitations, following the earlier suggestion of Refs. [13; 48]. The renormalization group analysis of this model with random uncorrelated interaction constants $u$ between each pair of defects [15; 40] predicted that the low energy excitations are distinguished by the number of participating defects $n$ and possess the universal densities of states $P_{n}\propto\ln(n)/n^{2}$. In that specific model the coupling constant for $n$-defect excitations scales with $n$ as $\gamma\propto\sqrt{n}$. The distribution of coupling constants $\gamma$ behaves consequently as
@@
The predicted distribution of dipole moments is relevant if the dipole-dipole interaction is the strongest long-range interaction in the system. To our knowledge in the vast majority of amorphous dielectrics elastic interaction is stronger than the dipole-dipole interaction [50; 16]. Indeed, TLSs of atomic nature possess very small dipole moments of order or less than few D because heavy atoms can tunnel only to very short distances. However, if TLSs are of electronic nature, then the tunneling distances can be much longer up to nanometers and the associated dipole-dipole interaction can exceed the elastic one. Consequently, we expect that the materials showing anomalous non-linear absorption possess TLSs of electronic nature formed by e. g. localized quasi-particles [51; 52].